DILUTED EARNING PER SHARE, CONVERTIBLE SECURITIES - ABC Corporation has weighted- average number of common stock outstanding throughout 2018 of 450,000 shares. In addition, the corporation has $6,000,000 bonds, 7%, convertible into 100,000 common stock. During the year 2018, the corporation has net income of $900,000. Required: Compute the basic earnings per share and the diluted earnings per share. The tax rate is 30%.
Answer #1

Earning per share = net income / number of outstanding shares = 900000/450000=$2

Diluted earning per share = (net income + interest after taxes) / outstanding shares

= (900000+(6000000*7%*(1-30%))/(450000+100000) = $2.17